# RE-RAG: Improving Open-Domain QA Performance and Interpretability with Relevance Estimator in Retrieval-Augmented Generation

### v2

@@ Abstract
The Retrieval Augmented Generation (RAG) framework utilizes a combination of parametric knowledge and external knowledge to demonstrate state-of-the-art performance on open-domain question answering tasks. However, the RAG framework suffers from performance degradation when the query is accompanied by irrelevant contexts. In this work, we propose the RE-RAG framework, which introduces a relevance estimator (RE) that not only provides relative relevance between contexts as previous rerankers did, but also provides confidence, which can be used to classify whether given context is useful for answering the given question. We propose a weakly supervised method for training the RE simply utilizing question-answer data without any labels for correct contexts. We show that RE trained with a small generator (sLM) can not only improve the sLM fine-tuned together with RE but also improve previously unreferenced large language models (LLMs). Furthermore, we investigate new decoding strategies that utilize the proposed confidence measured by RE such as choosing to let the user know that it is “unanswerable” to answer the question given the retrieved contexts or choosing to rely on LLM’s parametric knowledge rather than unrelated contexts.
@@ 1 Introduction
We demonstrate that RE-RAG, enhanced with RE, significantly improves upon the existing RAG. Addtionally, we show that RE trained on a small language model can improve the answer performance of LLMs. (§4.1)
@@ 2.3 Joint training of RE-RAG
| Model | Extra | Generator | NQ | TQA | # Contexts |
| Small language models () | |||||
| RAG (Lewis et al., 2020b) | - | 445M | 44.5 | 56.8 | 50 |
| (Izacard and Grave, 2021b) | - | 220M | 48.2 | 65.0 | 100 |
| (Izacard and Grave, 2021b) | - | 770M | 51.4 | 67.6 | 100 |
| (Izacard and Grave, 2021a) | - | 220M | 50.1 | 69.3 | 100 |
| (Izacard and Grave, 2021a) | - | 770M | 54.4 | 72.5 | 100 |
| ReAtt (Jiang et al., 2022) | - | 770M | 54.7 | - | 100 |
| (Izacard and Grave, 2021a) | - | 220M | 48.6 | 67.4 | 25 |
| (Izacard and Grave, 2021a) | - | 770M | 53.9 | 71.2 | 25 |
| R2-D2 (Fajcik et al., 2021) | 125M | 1.04B | 55.9 | 69.9 | 25 |
| 220M | 220M | 49.9 | 68.2 | 25 | |
| 770M | 220M | 51.4 | 69.5 | 25 | |
| 770M | 770M | 54.0 | 70.2 | 25 | |
| Large language models () | |||||
| (Asai et al., 2023) | - | 7B | - | 66.4 | 5 |
| (Asai et al., 2023) | - | 13B | - | 69.3 | 5 |
| 770M | 7B | 45.7 | 67.1 | 5 | |
| 770M | 13B | 46.6 | 70.8 | 5 | |
| RA-DIT (Lin et al., 2023) | - | 65B | 43.9 | 75.1 | 10 |
| - | 70B | 38.1(40.7) | 63.5(66.3) | 10 | |
| - | 70B | 39.5(46.8) | 68.1(72.1) | 10 | |
| - | 175B | 42.9(45.9) | 69.0(70.7) | 10 | |
| (Shi et al., 2023) | - | 175B | 45.5 | 77.3 | 10 |
| 770M | 8B | 49.6 | 73.0 | 10 | |
| 770M | 70B | 48.0 | 72.4 | 10 | |
| 770M | 70B | 50.8 | 75.5 | 10 | |
| 770M | 175B | 49.3 | 72.6 | 10 | |
@@ 3.3 Baseline
We consider an additional baseline to compare the performance of RE when applied to LLMs. We compare the performance of RE and FiD-KD retriever when applied to LLMs. When applying the FiD-KD retriever to LLMs, we compared two methods: traditional RAG, which uses the retriever similarity score to perform answer marginalization, and RALM, which concatenates all context. Furthermore, we compare our performance with other studies Asai et al. (2023); Lin et al. (2023); Shi et al. (2023) that have implemented RAG in LLMs.
@@ 3.4 Model
The two components of our framework, RE and the generator, utilize the T5 model (Raffel et al., 2020). We utilize the T5-base, T5-large models, and explore three different model sizes depending on the combination of the two models.
Additionally, we utilize Llama2 (7B, 13B, 70B), Llama3 111https://github.com/meta-llama/llama3 (8B, 70B), and ChatGPT (“gpt-3.5-turbo-0125” version) as generators to assess if RE brings performance improvements when applied to LLMs. In our experiments, the LLMs used as generators are not fine-tuned for the downstream task.
@@ 4.1 Main Results
The overall accuracy of our system on the two datasets (NQ and TQA) is shown in Table 1. Compared to the traditional RAG, our system, RE-RAG, performs better despite having the same total number of parameters. Our proposed RE improves the reliability of the RAG system by more accurately measuring the relevance between question and context. Our model performed competitively with models based on FiD structures(Izacard and Grave, 2021a; Jiang et al., 2022; Fajcik et al., 2021).

### v3

@@ Abstract
The Retrieval Augmented Generation (RAG) framework utilizes a combination of parametric knowledge and external knowledge to demonstrate state-of-the-art performance on open-domain question answering tasks. However, the RAG framework suffers from performance degradation when the query is accompanied by irrelevant contexts. In this work, we propose the RE-RAG framework, which introduces a relevance estimator (RE) that not only provides relative relevance between contexts as previous rerankers did, but also provides confidence, which can be used to classify whether given context is useful for answering the given question. We propose a weakly supervised method for training the RE simply utilizing question-answer data without any labels for correct contexts. We show that RE trained with a small generator (sLM) can not only improve the sLM fine-tuned together with RE but also improve previously unreferenced large language models (LLMs). Furthermore, we investigate new decoding strategies that utilize the proposed confidence measured by RE such as choosing to let the user know that it is “unanswerable” to answer the question given the retrieved contexts or choosing to rely on LLM’s parametric knowledge rather than unrelated contexts. 111Code is available at here
@@ 1 Introduction
We demonstrate that RE-RAG, enhanced with RE, significantly improves upon the existing RAG. Additionally, we show that RE trained on a small language model can improve the answer performance of LLMs. (§4.1)
@@ 2.3 Joint training of RE-RAG
| Model | Extra | Generator | NQ | TQA | # Contexts | ||||
| EM | Acc | F1 | EM | Acc | F1 | ||||
| Small language models () | |||||||||
| RAG (Lewis et al., 2020b) | - | 445M | 44.5 | - | - | 56.8 | - | - | 50 |
| (Izacard and Grave, 2021b) | - | 220M | 48.2 | - | - | 65.0 | - | - | 100 |
| (Izacard and Grave, 2021b) | - | 770M | 51.4 | - | - | 67.6 | - | - | 100 |
| (Izacard and Grave, 2021a) | - | 220M | 50.1 | - | - | 69.3 | - | - | 100 |
| (Izacard and Grave, 2021a) | - | 770M | 54.4 | - | - | 72.5 | - | - | 100 |
| ReAtt (Jiang et al., 2022) | - | 770M | 54.7 | - | - | - | - | - | 100 |
| (Izacard and Grave, 2021a) | - | 220M | 48.6 | - | - | 67.4 | - | - | 25 |
| (Izacard and Grave, 2021a) | - | 770M | 53.9 | - | - | 71.2 | - | - | 25 |
| R2-D2 (Fajcik et al., 2021) | 125M | 1.04B | 55.9 | - | - | 69.9 | - | - | 25 |
| 220M | 220M | 49.9 | 53.1 | 56.9 | 68.2 | 70.0 | 73.6 | 25 | |
| 220M | 220M | 51.9 | 55.2 | 58.9 | 70.1 | 72.0 | 75.8 | 25 | |
| 770M | 770M | 54.0 | 56.7 | 61.0 | 70.2 | 71.7 | 75.9 | 25 | |
| 770M | 770M | 55.4 | 58.3 | 62.5 | 72.9 | 74.4 | 78.7 | 25 | |
| Large language models () | |||||||||
| (Asai et al., 2023) | - | 7B | - | - | - | - | 66.4 | - | 5 |
| (Asai et al., 2023) | - | 13B | - | - | - | - | 69.3 | - | 5 |
| 770M | 7B | 45.7 | 48.4 | 54.3 | 67.1 | 70.1 | 73.3 | 5 | |
| 770M | 13B | 46.6 | 49.8 | 55.6 | 70.8 | 73.2 | 77.2 | 5 | |
| RA-DIT (Lin et al., 2023) | - | 65B | 43.9 | - | - | 75.1 | - | - | 10 |
| - | 8B | 37.9(38.2) | 43.9(40.2) | 47.5(47.4) | 63.8(57.6) | 66.7(59.3) | 70.7(63.3) | 10 | |
| - | 70B | 38.1(40.7) | 43.0(47.4) | 47.5(50.8) | 63.5(66.3) | 66.4(71.4) | 70.0(73.2) | 10 | |
| - | 70B | 39.5(46.8) | 44.3(52.4) | 48.5(56.9) | 68.1(72.1) | 70.8(75.3) | 74.7(79.1) | 10 | |
| - | 175B | 42.9(45.9) | 46.6(50.0) | 52.2(56.2) | 69.0(70.7) | 74.5(74.0) | 76.8(77.8) | 10 | |
| (Shi et al., 2023) | - | 175B | 45.5 | - | - | 77.3 | - | - | 10 |
| 770M | 8B | 49.6 | 54.5 | 59.0 | 73.0 | 75.4 | 79.3 | 10 | |
| 770M | 70B | 48.0 | 52.0 | 57.6 | 72.4 | 74.8 | 78.6 | 10 | |
| 770M | 70B | 50.8 | 54.8 | 60.1 | 75.5 | 77.7 | 81.7 | 10 | |
| 770M | 175B | 49.3 | 55.2 | 59.6 | 72.6 | 77.7 | 80.3 | 10 | |
@@ 3.2 Evaluation Metric
We also provide F1 score and accuracy (Acc) as an additional evaluation metric as some previous paper only report Acc (Asai et al., 2023), which assesses whether the generated string contains the gold answer. These scores show a similar trend with the EM score, that RE-RAG outperforms the baseline methods. Nonetheless, since most baselines report EM scores exclusively, our comparison is focused on EM scores.
@@ 3.3 Baseline
We consider an additional baseline to compare the performance of RE when applied to LLMs. We compare the performance of RE and FiD-KD retriever when applied to LLMs. When applying the FiD-KD retriever to LLMs, we compared two methods: traditional RAG, which uses the retriever similarity score to perform answer marginalization, and RALM, which concatenates all context. When generating answers for individual contexts using the traditional RAG method, we used 8-shot examples, while the RALM method employed a zero-shot approach due to context length limitations. Furthermore, we compare our performance with other studies Asai et al. (2023); Lin et al. (2023); Shi et al. (2023) that have implemented RAG in LLMs.
@@ 3.4 Model
The two components of our framework, RE and the generator, utilize the T5 model (Raffel et al., 2020) and Flan-T5 (Chung et al., 2024). We utilize the base and large size models, and explore three different model sizes depending on the combination of the two models.
Additionally, we utilize Llama2 (7B, 13B, 70B), Llama3 222https://github.com/meta-llama/llama3 (8B, 70B), and ChatGPT (“gpt-3.5-turbo-0125” version) as generators to assess if RE brings performance improvements when applied to LLMs. In our experiments, the LLMs used as generators are not fine-tuned for the downstream task.
@@ 4.1 Main Results
The overall accuracy of our system on the two datasets (NQ and TQA) is shown in Table 1. Compared to the traditional RAG, our system, RE-RAG, performs better despite having the same total number of parameters. Our proposed RE improves the reliability of the RAG system by more accurately measuring the relevance between question and context. Our model performed competitively with models based on FiD structures(Izacard and Grave, 2021a; Jiang et al., 2022; Fajcik et al., 2021). We also found that our methodology was more efficient than the instructed tuned T5.
@@ 5.1 Exploring decoding strategies in low confidence context sets
| Dataset | Model | Recall@k | |||
| R@1 | R@5 | R@10 | R@20 | ||
| NQ | FiD-KD(NQ NQ) | 49.4 | 73.8 | 79.6 | 84.3 |
| FiD-KD(TQA NQ) | 35.9 | 63.2 | 73.1 | 80.5 | |
| (NQ NQ) | 61.9 | 79.4 | 83.6 | 86.4 | |
| (TQA NQ) | 46.2 | 71.6 | 79.3 | 83.9 | |
| TQA | FiD-KD(TQA TQA) | 60.1 | 77.0 | 80.9 | 83.6 |
| FiD-KD(NQ TQA) | 47.6 | 70.8 | 76.8 | 81.1 | |
| (TQA TQA) | 70.4 | 82.2 | 84.4 | 86.1 | |
| (NQ TQA) | 67.8 | 80.2 | 83.0 | 85.1 | |
| Model | NQ (EM/Acc) | TQA (EM/Acc) | #Contexts |
| 37.9/43.9 | 63.8/66.7 | 10 | |
| 49.6/54.5 | 73.0/75.4 | 10 | |
| 30.3/34.7 | - | 10 | |
| 42.1/46.1 | - | 10 | |
| - | 57.6/60.4 | 10 | |
| - | 70.3/73.0 | 10 |
5.2 Evaluation of relevance estimator on unseen dataset
We evaluate the effectiveness of the relevance estimator (RE) module on unseen datasets that were not utilized during training from two perspectives: rerank performance and answer performance. The RE module and the baseline FiD-KD are trained only using a single dataset such as NQ (TQA). We analyze the changes in performance when applying the RE module and FiD-KD retriever to the new unseen dataset TQA (NQ).
Table 7 compares the rerank performance of the RE module and the FiD-KD retriever on datasets that were not referenced in training. Overall, the RE module consistently shows a smaller performance drop compared to the FiD-KD retriever on these unseen datasets. In particular, when the RE module trained on Natural Questions is extended to the TriviaQA dataset (NQTQA), both models show smaller performance drop than the opposite case (TQANQ). However, the performance drop of the RE module is notably smaller (-3.7%) than FiD-KD’s (-20.8%). This suggests that the RE module is more effective than FiD-KD retriever on unreferenced datasets when trained on datasets that are conducive to generalization.
Table 8 presents a comparison of the answer performance when the RE module and FiD-KD retriever are applied to the Llama3 8B generator on unreferenced datasets. The RE module consistently exhibits a smaller performance degradation compared to the FiD-KD retriever on both datasets, similar to the recall performance in Table 7.
@@ 8 Limitation
9 Acknowledgements
This work was supported in part by the National Research Foundation of Korea (NRF) grant (RS-2023-00280883, RS-2023-00222663) and New Faculty Startup Fund from Seoul National University, and with the aid of computing resources from Artificial Intelligence Industry Center Agency, and Google cloud platform research credits, and National Super computing Center with super computing resources including technical support (KSC-2023-CRE-0176).